# Fisher-Hartwig asymptotics for non-Hermitian random matrices

### v1

@@ 3 Singular potentials
$\times\prod_{j=1}^{m}e^{\frac{\gamma_{j}}{2}\big(\int_{\partial S}f\sigma{\rm d}s-f(\zeta_{j})\big)}\prod_{j=1}^{m}e^{\frac{\gamma_{j}^{2}}{8\pi}\int_{\mathbb{C}}\nabla(\log_{\delta}^{\zeta_{j}})^{S}\cdot\nabla(2\log_{\Delta}^{\zeta_{j}}-\log_{\delta}^{\zeta_{j}})^{S}\,{\rm d}m+\frac{\gamma_{j}}{4}(1+L(\zeta_{j})-L^{S}_{\infty})}$
$\times\mathbb{E}\Big[\prod_{j=1}^{m}e^{\sum_{i=1}^{N}\gamma_{j}(\log_{\Delta}^{\zeta_{j}}-\log_{\delta}^{\zeta_{j}})(z_{i})}\Big]\prod_{j\neq k}|\zeta_{j}-\zeta_{k}|^{-\frac{\gamma_{i}\gamma_{j}}{4}}e^{\frac{\gamma_{i}\gamma_{j}}{4}\int_{\partial S}\log^{\zeta_{j}}\sigma{\rm d}s}\left(1+\operatorname{O}(N^{-\kappa/10})\right).$
@@
We begin similarly to the proof of Proposition 2.7, with the exception that we remove only the function $f$ and the mesoscopic contributions from logarithmic singularities using Proposition 2.6. More explicitly, let $\chi$ be an order $1$ bump function that is equal to $1$ on $S\cup S_{\rm edge}(1/2)$ and 0 on $S_{\rm out}(2/3)$. Then $\sum_{j}\gamma_{j}(\log_{\Delta}^{\zeta_{j}}-\log_{\delta}^{\zeta_{j}})$ and $f+\sum_{j}\gamma_{j}\log_{\delta}^{\zeta_{j}}$ substituted into $f$ and $g$ respectively (with $g_{1}=g\chi$, $g_{2}=g(1-\chi)$) in Proposition 2.6 yields
@@
$\Gamma(\log^{\zeta_{j}})$ $=\frac{1}{8\pi}\int_{S}\Delta\log_{\delta}^{\zeta_{j}}\,{\rm d}m+\frac{1}{8\pi}\int_{S}\log_{\delta}^{\zeta_{j}}\Delta L\,{\rm d}m-\frac{1}{8\pi}\Big(\int_{\partial S}\log_{\delta}^{\zeta_{j}}\mathcal{N}(L){\rm d}s\Big),$
@@
We now evaluate each term one-by-one. $\Gamma(f)$ is already in the desired form. Recalling the definition of the log-jump measure $\sigma$ given in (1.4) and the fact that $\Delta(\log^{\zeta}_{\Delta})=2\pi\frac{\chi^{\zeta}_{\Delta}}{\|\chi^{\zeta}_{\Delta}\|_{L^{1}}}$, $\Gamma(f,\log^{\zeta_{j}})$ can be simplified easily as follows:
@@
$=-\frac{1}{2}f(\zeta_{j})+\frac{1}{2}\int_{\partial S}f\,\sigma\,{\rm d}s+\operatorname{O}(N^{-\kappa}).$
Next, for $\Gamma(\log^{\zeta_{j}})$ term, we begin by establishing the analogue of (2.17), accounting for the non-compact support of the logarithm. We modify $L^{S}$ to $L^{S}-L^{S}_{\infty}$ so that Green’s identity can be applied on $S^{\rm c}$:
@@
$=-\frac{1}{4}\log|\zeta_{j}-\zeta_{k}|+\frac{1}{4}\int_{\partial S}\log^{\zeta_{j}}\sigma\,{\rm d}s+\operatorname{O}(N^{-\kappa}).$
Therefore it remains only to show that $\mathscr{E}_{N}(f+tg,g_{1})=\operatorname{O}(N^{-\kappa/10})$. Let $h_{t}=f+tg$. Since the expression is linear in its second argument, it suffices to establish this bound for $\mathscr{E}_{N}(h_{t},g)$ uniformly on $g\in\mathscr{A}_{C,\tau}$ with $\tau\in[N^{-1/2+\kappa},1]$, which is done in the next lemma. ∎
@@
The proof follows the same general strategy as in Lemmas 2.82.10, with only a few modifications that we now indicate. Most of the steps in the proofs of the lemmas remain valid without change: in particular, using the notation from Appendix A, we still have $\widetilde{\bf K}_{N}=\operatorname{O}(N)$, exponential decay of $\widetilde{\bf K}_{N}(z,z)$ outside the unit disc, and $|\frac{\widetilde{\bf K}_{N}(z,z)}{N}-\rho_{V}(z)|\leqslant N^{-2\kappa+\varepsilon}$ when $z\in S_{\rm in}(\delta_{N})\setminus\cup_{j}{\rm B}(\zeta_{j},\delta)$ by Lemma A.4 and Theorem A.1. The stability of the overwhelming probability events also continues to hold by Corollary 3.9. The only points where the argument requires modification are those that invoke integration in a neighbourhood of the singularities, i.e. on $\cup_{j}{\rm B}(\zeta_{j},\delta)$. In these steps, the local behaviour around the singularities must be handled separately, as we describe below.
@@
$\left|\mathbb{E}_{f}\Big[\int_{S_{\rm in}(\delta_{N})\cap{\rm B}(\zeta_{j},3\delta_{N})}\frac{\bar{\partial}p}{\Delta V}\partial f{\rm d}\tilde{\mu}_{N}\Big]\right|\lesssim\int_{{\rm B}(\zeta_{j},3\delta_{N})}\frac{1}{\tau}\frac{1}{|z-\zeta_{j}|}{\rm d}m(z)\lesssim\frac{\delta_{N}}{\tau}=\operatorname{O}(N^{-\kappa+\varepsilon})$ (3.1)
@@
$\left|\mathbb{E}_{f}\Big[\int_{S_{\rm in}(\delta_{N})\cap{\rm B}(\zeta_{j},\delta)\setminus{\rm B}(\zeta_{j},3\delta_{N})}\frac{\bar{\partial}p}{\Delta V}\partial f{\rm d}\tilde{\mu}_{N}\Big]\right|\lesssim\int_{{\rm B}(\zeta_{j},\delta)\setminus{\rm B}(\zeta_{j},3\delta_{N})}\frac{1}{\tau}\frac{1}{|z-\zeta_{j}|}\frac{1}{N|z-\zeta_{j}|^{2}}{\rm d}m(z)\lesssim\frac{1}{\tau\delta_{N}N}=\operatorname{O}(N^{-\kappa})$ (3.2)
@@
$E_{f}\Big[\int_{{\rm B}(\zeta_{j},3\delta_{N})}\frac{\Delta g}{\Delta V}\,{\rm d}\tilde{\mu}_{N}\Big]=\operatorname{O}(\frac{\delta_{N}^{2}}{\tau^{2}}),\quad E_{f}\Big[\int_{S_{\rm in}(\delta_{N})\cap{\rm B}(\zeta_{j},\delta)\setminus{\rm B}(\zeta_{j},3\delta_{N})}\frac{\Delta g}{\Delta V}\,{\rm d}\tilde{\mu}_{N}\Big]=\operatorname{O}(\frac{\log N}{N\tau^{2}}).$
@@
$E_{f}\Big[\int_{{\rm B}(\zeta_{j},\delta)}\frac{\bar{\partial}p\partial L}{\Delta V/4}\,{\rm d}\tilde{\mu}_{N}\Big]=\operatorname{O}(\delta^{2}/\tau)$
@@
$\mathbb{E}_{f}\Big[N\iint_{z\neq w}\frac{r(z)-r(w)}{z-w}\tilde{\mu}_{N}({\rm d}z)\tilde{\mu}_{N}({\rm d}w)\Big]=\operatorname{O}(N^{-\kappa+\varepsilon}),$ (3.3)
@@
$\mathbb{E}\Big[e^{\sum_{i=1}^{N}f(z_{i})}\prod_{j=1}^{m}e^{\sum_{i=1}^{N}\gamma_{j}\log_{\Delta}(z_{i}-\zeta_{j})}\Big]=e^{N\int f+\sum_{j}\gamma_{j}\log_{\delta}^{\zeta_{j}}{\rm d}\mu_{V}}e^{\frac{1}{8\pi}\big(\int_{\mathbb{C}}|\nabla f^{S}|^{2}{\rm d}m+\mathds{1}_{S}\Delta f+L^{S}\Delta f{\rm d}m\big)}\\ \times\prod_{j=1}^{m}e^{\frac{\gamma_{j}}{2}\big(\int_{\partial S}f\sigma{\rm d}s-f(\zeta_{j})\big)}\prod_{j=1}^{m}e^{\frac{\gamma_{j}^{2}}{8\pi}\int_{\mathbb{C}}\nabla(\log_{\delta}^{\zeta_{j}})^{S}\cdot\nabla(2\log_{\Delta}^{\zeta_{j}}-\log_{\delta}^{\zeta_{j}})^{S}\,{\rm d}m+\frac{\gamma_{j}}{4}(1+L(\zeta_{j})-L^{S}_{\infty})}\\ \times\prod_{j=1}^{m}\mathbb{E}^{\rm Gin}\left[e^{\sum_{i=1}^{N}\gamma_{j}(\log_{\Delta_{j}}-\log_{\delta_{j}})(z_{i})}\right]\prod_{j\neq k}|\zeta_{j}-\zeta_{k}|^{-\frac{\gamma_{i}\gamma_{j}}{4}}e^{\frac{\gamma_{i}\gamma_{j}}{4}\int_{\partial S}\log^{\zeta_{j}}\sigma{\rm d}s}\left(1+\operatorname{O}(N^{-\kappa/10})\right)$
@@
$\mathbb{E}^{\rm Gin}\Big[\prod_{i=1}^{N}|z_{i}|^{\gamma_{j}}\Big]=$ $\,e^{\frac{N}{\pi}\int_{\mathbb{D}}\log_{\delta_{j}}{\rm d}m}e^{\gamma_{j}^{2}\int_{\mathbb{C}}\nabla(\log_{\delta_{j}})^{\mathbb{D}}\cdot\nabla(2\log_{\Delta_{j}}-\log_{\delta_{j}})^{\mathbb{D}}\,{\rm d}m+\frac{\gamma_{j}}{4}}$
@@
$=\pi L(\zeta_{j})+2\pi\int_{\partial S}\log^{\zeta_{j}}\sigma\,{\rm d}s+\operatorname{O}(N^{-\kappa-\alpha}).$
@@ 4 Convergence to the Gaussian multiplicative chaos
where $\mathfrak{s}$ is as in Definition 1.1. Then
@@ Appendix A Kernel estimates
Appendix B The log-jump measure
Here we prove the claims that are made in Definition 1.1.
Well-definedness of (1.4). We first show that the function $\mathcal{N}(\log^{\zeta})$ does not depend on the choice of $\zeta\in S$.
@@ Appendix B The log-jump measure
$\sigma$ is non-negative. Fix an arbitrary point $\zeta\in S$. Let $f=\log^{\zeta}-(\log^{\zeta})^{S}$ on $S^{\rm c}$. Note that $f$ is harmonic on $S^{\rm c}$, $f=0$ on $\partial S$ and $f(z)$ goes to $\infty$ as $|z|\to\infty$. By the maximum principle, $f$ attains its minimum value on $\partial S$. Thus, $f(z)\geqslant 0$ for all $z\in S^{\rm c}$. Thus, $\frac{\partial f}{\partial n}\geqslant 0$ on $\partial S$.
$\sigma$ is a probability measure. What remains is to verify that $\int_{\partial S}\sigma\,{\rm d}s$ is $1$. This follows from the Green’s theorem,
@@
$\mathfrak{s}$ is well-defined. It suffices to show that $\int_{\partial S}\log^{\zeta_{1}}\sigma\,{\rm d}s=\int_{\partial S}\log^{\zeta_{2}}\sigma\,{\rm d}s$ for any $\zeta_{1},\zeta_{2}$ distinct in the interior of $S$. We begin with a symmetric expression, and apply Green’s theorem again,
$\int_{\mathbb{C}}\nabla(\log^{\zeta_{1}})^{S}\cdot\nabla(\log^{\zeta_{2}})^{S}{\rm d}m=\int_{S}\nabla(\log^{\zeta_{1}})^{S}\cdot\nabla(\log^{\zeta_{2}})^{S}{\rm d}m+\int_{S^{\rm c}}\nabla(\log^{\zeta_{1}})^{S}\cdot\nabla(\log^{\zeta_{2}})^{S}{\rm d}m\\ =\int_{\partial S}\log^{\zeta_{1}}\frac{\partial\log^{\zeta_{2}}}{\partial n}{\rm d}s-\int_{S}\log^{\zeta_{1}}\Delta\log^{\zeta_{2}}{\rm d}m+\int_{\partial S}\log^{\zeta_{1}}\frac{\partial(\log^{\zeta_{2}})^{S}}{\partial(-n)}{\rm d}s=-2\pi\log|\zeta_{1}-\zeta_{2}|+2\pi\int_{\partial S}\log^{\zeta_{1}}\sigma\,{\rm d}s.$
@@
$\int_{\partial S}\log^{\zeta_{1}}\sigma\,{\rm d}s=\int_{\partial S}\log^{\zeta_{2}}\sigma\,{\rm d}s.$
@@
The properties of the log-jump measure stated in Definition 1.1 and proved above can easily be extended to the case of arbitrary connected droplet $S$ with smooth boundary, without the simple connectivity assumption. In this general case, independence of $\sigma$ from the choice of $\zeta$ is proved in the same manner, $\sigma$ is a measure with mass $1$ which is nonnegative and supported on the outer boundary of $S$, and $\int_{\partial S}\log^{\zeta}\sigma\,{\rm d}s$ does not depend on $\zeta$, again.

### v2

@@ 3 Singular potentials
$\times\prod_{j=1}^{m}e^{\frac{\gamma_{j}}{2}\big(\int_{\partial S}f{\rm d}\omega^{\infty}-f(\zeta_{j})\big)}\prod_{j=1}^{m}e^{\frac{\gamma_{j}^{2}}{8\pi}\int_{\mathbb{C}}\nabla(\log_{\delta}^{\zeta_{j}})^{S}\cdot\nabla(2\log_{\Delta}^{\zeta_{j}}-\log_{\delta}^{\zeta_{j}})^{S}\,{\rm d}m+\frac{\gamma_{j}}{4}(1+L(\zeta_{j})-L^{S}(\infty))}$
$\times\mathbb{E}\Big[\prod_{j=1}^{m}e^{\sum_{i=1}^{N}\gamma_{j}(\log_{\Delta}^{\zeta_{j}}-\log_{\delta}^{\zeta_{j}})(z_{i})}\Big]\prod_{j\neq k}|\zeta_{j}-\zeta_{k}|^{-\frac{\gamma_{i}\gamma_{j}}{4}}e^{\frac{\gamma_{i}\gamma_{j}}{4}\mathfrak{s}}\left(1+\operatorname{O}(N^{-\kappa/10})\right).$
@@
We begin similarly to the proof of Proposition 2.7, with the exception that we remove only the function $f$ and the mesoscopic contributions from logarithmic singularities using Proposition 2.6. More explicitly, let $\chi$ be an order $1$ bump function that is equal to $1$ on $S\cup S_{\rm edge}(1/2)$ and 0 on $S_{\rm out}(2/3)$. Then $F:=\sum_{j}\gamma_{j}(\log_{\Delta}^{\zeta_{j}}-\log_{\delta}^{\zeta_{j}})$ and $g:=f+\sum_{j}\gamma_{j}\log_{\delta}^{\zeta_{j}}$ substituted into $f$ and $g$ respectively (with $g_{1}=g\chi$, $g_{2}=g(1-\chi)$) in Proposition 2.6 yields
@@
$\Gamma(\log^{\zeta_{j}})$ $=\frac{1}{8\pi}\int_{S}\Delta\log_{\delta}^{\zeta_{j}}\,{\rm d}m+\frac{1}{8\pi}\int_{S}\log_{\delta}^{\zeta_{j}}\Delta L\,{\rm d}m-\frac{1}{8\pi}\int_{\partial S}\log_{\delta}^{\zeta_{j}}\mathcal{N}(L){\rm d}s,$
@@
We now evaluate each term one-by-one. $\Gamma(f)$ is already in the desired form. Recalling the definition of the harmonic measure $\omega^{\infty}$ given in (1.6) and the fact that $\Delta(\log^{\zeta}_{\Delta})=2\pi\frac{\chi^{\zeta}_{\Delta}}{\|\chi^{\zeta}_{\Delta}\|_{L^{1}}}$, $\Gamma(f,\log^{\zeta_{j}})$ can be simplified easily as follows:
@@
$=-\frac{1}{2}f(\zeta_{j})+\frac{1}{2}\int_{\partial S}f\,{\rm d}\omega^{\infty}+\operatorname{O}(N^{-\kappa}).$
Next, for $\Gamma(\log^{\zeta_{j}})$ term, we begin by establishing the analogue of (2.17), accounting for the non-compact support of the logarithm. We denote $L^{S}(\infty)$ by $L^{S}_{\infty}$ and modify $L^{S}$ to $L^{S}-L^{S}_{\infty}$ so that Green’s identity can be applied on $S^{\rm c}$:
@@
$=-\frac{1}{4}\log|\zeta_{j}-\zeta_{k}|+\frac{1}{4}\int_{\partial S}\log^{\zeta_{j}}{\rm d}\omega^{\infty}+\operatorname{O}(N^{-\kappa})$
and the second integral in the right hand side is equal to $\mathfrak{s}$ by (1.7). Therefore it remains only to show that $\mathscr{E}_{N}(F+tg,g_{1})=\operatorname{O}(N^{-\kappa/10})$. Let $h_{t}=F+tg$. Since the expression is linear in its second argument, it suffices to establish this bound for $\mathscr{E}_{N}(h_{t},g)$ uniformly on $g\in\mathscr{A}_{C,\tau}$ with $\tau\in[N^{-1/2+\kappa},1]$, which is done in the next lemma. ∎
@@
The proof follows the same general strategy as in Lemmas 2.8-2.10, with only a few modifications that we now indicate. Most of the steps in the proofs of the lemmas remain valid without change: in particular, using the notation from Appendix A, we still have $\widetilde{\bf K}_{N}=\operatorname{O}(N)$, exponential decay of $\widetilde{\bf K}_{N}(z,z)$ outside the unit disc, and $|\frac{\widetilde{\bf K}_{N}(z,z)}{N}-\rho_{V}(z)|\leqslant N^{-2\kappa+\varepsilon}$ when $z\in S_{\rm in}(\delta_{N})\setminus\cup_{j}{\rm B}(\zeta_{j},\delta)$ by Lemma A.4 and Theorem A.1. The stability of the overwhelming probability events also continues to hold by Corollary 3.9. The only points where the argument requires modification are those that invoke integration in a neighbourhood of the singularities, i.e. on $\cup_{j}{\rm B}(\zeta_{j},\delta)$. In these steps, the local behaviour around the singularities must be handled separately, as we describe below.
@@
$\left|\mathbb{E}_{h_{t}}\Big[\int_{S_{\rm in}(\delta_{N})\cap{\rm B}(\zeta_{j},3\delta_{N})}\frac{\bar{\partial}p}{\Delta V}\partial f{\rm d}\tilde{\mu}_{N}\Big]\right|\lesssim\int_{{\rm B}(\zeta_{j},3\delta_{N})}\frac{1}{\tau}\frac{1}{|z-\zeta_{j}|}{\rm d}m(z)\lesssim\frac{\delta_{N}}{\tau}=\operatorname{O}(N^{-\kappa+\varepsilon})$ (3.1)
@@
$\left|\mathbb{E}_{h_{t}}\Big[\int_{S_{\rm in}(\delta_{N})\cap{\rm B}(\zeta_{j},\delta)\setminus{\rm B}(\zeta_{j},3\delta_{N})}\frac{\bar{\partial}p}{\Delta V}\partial f{\rm d}\tilde{\mu}_{N}\Big]\right|\lesssim\int_{{\rm B}(\zeta_{j},\delta)\setminus{\rm B}(\zeta_{j},3\delta_{N})}\frac{1}{\tau}\frac{1}{|z-\zeta_{j}|}\frac{1}{N|z-\zeta_{j}|^{2}}{\rm d}m(z)\lesssim\frac{1}{\tau\delta_{N}N}=\operatorname{O}(N^{-\kappa})$ (3.2)
@@
$E_{h_{t}}\Big[\int_{{\rm B}(\zeta_{j},3\delta_{N})}\frac{\Delta g}{\Delta V}\,{\rm d}\tilde{\mu}_{N}\Big]=\operatorname{O}(\frac{\delta_{N}^{2}}{\tau^{2}}),\quad E_{h_{t}}\Big[\int_{S_{\rm in}(\delta_{N})\cap{\rm B}(\zeta_{j},\delta)\setminus{\rm B}(\zeta_{j},3\delta_{N})}\frac{\Delta g}{\Delta V}\,{\rm d}\tilde{\mu}_{N}\Big]=\operatorname{O}(\frac{\log N}{N\tau^{2}}).$
@@
$E_{h_{t}}\Big[\int_{{\rm B}(\zeta_{j},\delta)}\frac{\bar{\partial}p\partial L}{\Delta V/4}\,{\rm d}\tilde{\mu}_{N}\Big]=\operatorname{O}(\delta^{2}/\tau)$
@@
$\mathbb{E}_{h_{t}}\Big[N\iint_{z\neq w}\frac{r(z)-r(w)}{z-w}\tilde{\mu}_{N}({\rm d}z)\tilde{\mu}_{N}({\rm d}w)\Big]=\operatorname{O}(N^{-\kappa+\varepsilon}),$ (3.3)
@@
$\mathbb{E}\Big[e^{\sum_{i=1}^{N}f(z_{i})}\prod_{j=1}^{m}e^{\sum_{i=1}^{N}\gamma_{j}\log_{\Delta}(z_{i}-\zeta_{j})}\Big]=e^{N\int f+\sum_{j}\gamma_{j}\log_{\delta}^{\zeta_{j}}{\rm d}\mu_{V}}e^{\frac{1}{8\pi}\big(\int_{\mathbb{C}}|\nabla f^{S}|^{2}{\rm d}m+\mathds{1}_{S}\Delta f+L^{S}\Delta f{\rm d}m\big)}\\ \times\prod_{j=1}^{m}e^{\frac{\gamma_{j}}{2}\big(\int_{\partial S}f{\rm d}\omega^{\infty}-f(\zeta_{j})\big)}\prod_{j=1}^{m}e^{\frac{\gamma_{j}^{2}}{8\pi}\int_{\mathbb{C}}\nabla(\log_{\delta}^{\zeta_{j}})^{S}\cdot\nabla(2\log_{\Delta}^{\zeta_{j}}-\log_{\delta}^{\zeta_{j}})^{S}\,{\rm d}m+\frac{\gamma_{j}}{4}(1+L(\zeta_{j})-L^{S}(\infty))}\\ \times\prod_{j=1}^{m}\mathbb{E}^{\rm Gin}\left[e^{\sum_{i=1}^{N}\gamma_{j}(\log_{\Delta_{j}}-\log_{\delta_{j}})(z_{i})}\right]\prod_{j\neq k}|\zeta_{j}-\zeta_{k}|^{-\frac{\gamma_{i}\gamma_{j}}{4}}e^{\frac{\gamma_{i}\gamma_{j}}{4}\mathfrak{s}}\left(1+\operatorname{O}(N^{-\kappa/10})\right)$
@@
$\mathbb{E}^{\rm Gin}\Big[\prod_{i=1}^{N}|z_{i}|^{\gamma_{j}}\Big]=$ $\,e^{\frac{N}{\pi}\gamma_{j}\int_{\mathbb{D}}\log_{\delta_{j}}{\rm d}m}e^{\gamma_{j}^{2}\int_{\mathbb{C}}\nabla(\log_{\delta_{j}})^{\mathbb{D}}\cdot\nabla(2\log_{\Delta_{j}}-\log_{\delta_{j}})^{\mathbb{D}}\,{\rm d}m+\frac{\gamma_{j}}{4}}$
@@
$=\pi L(\zeta_{j})+2\pi\mathfrak{s}+\operatorname{O}(N^{-\kappa-\alpha}).$
@@ 4 Convergence to the Gaussian multiplicative chaos
where $\mathfrak{s}$ is as in (1.4). Then
@@ Appendix A Kernel estimates
Appendix B The harmonic measure and capacity
In this appendix, we justify the definitions of the harmonic measure seen from infinity (1.6) and the logarithmic capacity (1.7), and prove several basic properties mentioned in the introduction.
Well-definedness of (1.6). We show that the function $\mathcal{N}(\log^{\zeta})$ does not depend on the choice of $\zeta\in S$.
@@ Appendix B The harmonic measure and capacity
$\omega^{\infty}$ is non-negative. Fix an arbitrary point $\zeta\in S$. Let $f=\log^{\zeta}-(\log^{\zeta})^{S}$ on $S^{\rm c}$. Note that $f$ is harmonic on $S^{\rm c}$, $f=0$ on $\partial S$ and $f(z)$ goes to $\infty$ as $|z|\to\infty$. By the maximum principle, $f$ attains its minimum value on $\partial S$. Thus, $f(z)\geqslant 0$ for all $z\in S^{\rm c}$. Thus, $\frac{\partial f}{\partial n}\geqslant 0$ on $\partial S$.
$\omega^{\infty}$ is a probability measure. What remains is to verify that $\int_{\partial S}{\rm d}\omega^{\infty}$ is $1$. This follows from the Green’s theorem,
@@
Well-definedness of (1.4). By the second equality in (1.7), it suffices to show that $\int_{\partial S}\log^{\zeta_{1}}{\rm d}\omega^{\infty}=\int_{\partial S}\log^{\zeta_{2}}{\rm d}\omega^{\infty}$ for any $\zeta_{1},\zeta_{2}$ distinct in the interior of $S$. We begin with a symmetric expression, and apply Green’s theorem again,
$\int_{\mathbb{C}}\nabla(\log^{\zeta_{1}})^{S}\cdot\nabla(\log^{\zeta_{2}})^{S}{\rm d}m=\int_{S}\nabla(\log^{\zeta_{1}})^{S}\cdot\nabla(\log^{\zeta_{2}})^{S}{\rm d}m+\int_{S^{\rm c}}\nabla(\log^{\zeta_{1}})^{S}\cdot\nabla(\log^{\zeta_{2}})^{S}{\rm d}m\\ =\int_{\partial S}\log^{\zeta_{1}}\frac{\partial\log^{\zeta_{2}}}{\partial n}{\rm d}s-\int_{S}\log^{\zeta_{1}}\Delta\log^{\zeta_{2}}{\rm d}m+\int_{\partial S}\log^{\zeta_{1}}\frac{\partial(\log^{\zeta_{2}})^{S}}{\partial(-n)}{\rm d}s=-2\pi\log|\zeta_{1}-\zeta_{2}|+2\pi\int_{\partial S}\log^{\zeta_{1}}{\rm d}\omega^{\infty}.$
@@
$\int_{\partial S}\log^{\zeta_{1}}{\rm d}\omega^{\infty}=\int_{\partial S}\log^{\zeta_{2}}{\rm d}\omega^{\infty}.$
@@
The properties of the harmonic measure stated in the introduction and proved above can easily be extended to the case of arbitrary connected droplet $S$ with smooth boundary, without the simple connectivity assumption. In this general case, independence of $\omega^{\infty}$ from the choice of $\zeta$ is proved in the same manner, $\omega^{\infty}$ is a measure with mass $1$ which is non-negative and supported on the outer boundary of $S$, and $\int_{\partial S}\log^{\zeta}{\rm d}\omega^{\infty}$ does not depend on $\zeta$, again.